OpenStudy (anonymous):

Solve sin^2x = cos^2x

OpenStudy (jhannybean):

hint : sin^2 (x) = 1- cos^2 x

OpenStudy (jhannybean):

and cos^2 (x) + sin^2 (x) = 1 ANOTHER HINT!!

OpenStudy (jhannybean):

OMG so many hints :)

Parth (parthkohli):

The 2nd is just a rip-off of the first. :-P

OpenStudy (anonymous):

Hints don't help when you have no idea what you're doing O.o

OpenStudy (jhannybean):

lolol.

OpenStudy (luigi0210):

Use your identities.

OpenStudy (jhannybean):

well!!! you need to know your identities mr.!! :P

OpenStudy (luigi0210):

can I be "Isuckatmath9998"?

OpenStudy (anonymous):

It's not even true, I'm pretty good at math, I just have a horrible teacher xD

Parth (parthkohli):

I'd be "isuckatmath10000000" because I suck at it a little too much.

OpenStudy (jhannybean):

noooo don't blame the teacher xD

OpenStudy (anonymous):

It's an online course, we literally have like nothing to reference

OpenStudy (luigi0210):

there are some horrible teachers out there :/

OpenStudy (jhannybean):

OK you have sin^2 (x) + cos^2 (x) = 1 RIGHT??

Parth (parthkohli):

@isuckatmath9999 You have the whole internet.

OpenStudy (zzr0ck3r):

1/2= cos(x)^2 can you do that?

OpenStudy (luigi0210):

Here I'll get you started: \[\sin^2x=(1-\sin^2x)\]

OpenStudy (jhannybean):

WTH IS 1/2 = COS ^ (X)?

OpenStudy (luigi0210):

zz made a booboo

OpenStudy (jhannybean):

lol??? tell me more about this identity. :)

OpenStudy (anonymous):

I feel like this is counterproductive

OpenStudy (luigi0210):

wait..

OpenStudy (jhannybean):

zz pulls identities out of his butt.

Parth (parthkohli):

\[\sin^2(x) = \dfrac{1}{2} \Rightarrow \sin(x) = \pm \frac{1}{\sqrt{2}}\]

OpenStudy (jhannybean):

can you elaborate....because im confused now. :|

OpenStudy (zzr0ck3r):

sin(x)^2 = cos(x)^2 imples cos(x)^2 = 1-cos(x)^2 imples 2cos(x)^2 = 1 imples cos(x)^2 = 1/2

OpenStudy (luigi0210):

there are two ways to solve this

Parth (parthkohli):

Should be 2 solutions.

OpenStudy (luigi0210):

do we have any restrictions?

OpenStudy (zzr0ck3r):

infinite solutions

OpenStudy (zzr0ck3r):

@Luigi0210 what I do wrong?

OpenStudy (anonymous):

0 ≤ x ≤ 2π

Parth (parthkohli):

Two different solutions in the interval \(0 \le x < 2\pi\)

sam (.sam.):

Ugh is it \[\sin^2x = \cos^2x\] why not just \[\tan^2x=1\]? then \[tanx= \pm 1\]

OpenStudy (zzr0ck3r):

we are all saying the same thing......

OpenStudy (luigi0210):

Nothing, sorry about that zz ^_^'

OpenStudy (luigi0210):

ha, math is fun .-.

OpenStudy (zzr0ck3r):

x = (pi)n/2-pi/4

OpenStudy (zzr0ck3r):

pick n according to restrictions...

sam (.sam.):



OpenStudy (jhannybean):

Let'sall stick with @Mr. Kohli's and @zzr0ck3r 's methods.....

OpenStudy (zzr0ck3r):

I think @.sam was the best

OpenStudy (zzr0ck3r):

@.Sam.

Parth (parthkohli):

Indeed, I didn't think of what .Sam. said :P

sam (.sam.):

There will be 4 solutions x=45 x=180-45=135 x=180+45=225 x=360-45=315

OpenStudy (jhannybean):

Tbh i havent even seen what Sam said, looking now .lol.

Parth (parthkohli):

Oh

OpenStudy (jhannybean):

Oh wow. I see. But does @isuckatmath9999 understand @.Sam.'s method?

OpenStudy (anonymous):

Yes i do that really helped :)

sam (.sam.):

\[\sin^2x=\cos^2x \] Divide both sides by \(cos^2x\) \[\frac{\sin^2x}{\cos^2x}=1\] Using \[tanx=\frac{sinx}{cosx}\] So \[\tan^2x=1\] Square roo both sides \[\tan(x)= \pm 1\] Then use the the diagram (forgot what it's called) for tangent  Since it's \(\pm 1\) you'll have 4 solutions, but sometimes it doesnt, you'll have to be careful for that

OpenStudy (jhannybean):

Awesome